KyledaGreat:

Evaluate this function at x=−7. Express your answer as an integer or simplified fraction. If the function is undefined at the given value, indicate "Undefined".

KyledaGreat:

1 attachment
KyledaGreat:

@lowkey

lowkey:

What do you think?

KyledaGreat:

\[\frac{ 3 }{ 2}\]

lowkey:

Let me do my magic and check if it's right

KyledaGreat:

okay

lowkey:

Not quite. How did you get to this answer?

KyledaGreat:

\[\frac{ 1 }{ 2} (-7) + 5\]

lowkey:

Okay. So restrictions are basically the ones with the greater sign signs and the x. You follow?

KyledaGreat:

I'm behind you so far

KyledaGreat:

it's not the right answer ?

KyledaGreat:

low ?

lowkey:

hellow! are you online? sorry i was afk.

lowkey:

So restrictions basically for the first equation is saying, for any number less than -3, use this formula. if your number is greater than or equal to -3, use the second formula. Now that we know this, which formula do we use to plug in -7?

KyledaGreat:

yes

KyledaGreat:

the first equation ?

KyledaGreat:

and it's alright

lowkey:

Yes! so plug in -7 to first equation

KyledaGreat:

\[-\frac{ 3 }{ 4}(-7) - 7 = -\frac{ 7 }{ 4}\]